in a sale,a chlothes shop reduces its prices by 25%. a shirt usally cost 36 pounds. how much is it in the sale?

Answer by Theo(13342)

retail price is 36 pounds
discount is 25%.
.25 * 36 = 9
discount price = 36 - 9 = 27.

another way to calculate it:
retail price is 36.
discount is 25% = .25
discount price is 36 - .25 * 36 = (1 - .25) * 36 = .75 * 36 = 27